Solve the following equation using inverse method

[1 3 0] [x] [4]

[0 0.5 1] [y]=[1]

[0.5 0 1] [z] [3]


1
Expert's answer
2022-03-15T19:23:55-0400

Solution


Given that


"\begin{bmatrix}
 1 & 3 & 0\\
 0 & 0.5 & 1\\
 0.5 & 0&1\\
\end{bmatrix}\times\begin{bmatrix}
 x \\
 y \\
 z\\
\end{bmatrix}=\begin{bmatrix}
 4 \\
 1 \\
3\\
\end{bmatrix}"


This can be written as


"AX=B"


"A^{-1}AX=A^{-1}B"


"IX=A^{-1}B"


"X=A^{-1}B"


Now


"A=\begin{bmatrix}
 1 & 3 & 0\\
 0 & 0.5 & 1\\
 0.5 & 0&1\\
\end{bmatrix}"


"\begin{gathered}
 \Rightarrow & \det \left( A \right) = 1\left( {0.5 - 0} \right) - 3\left( {0 - 0.5} \right) + 0\left( {0 - 0.25} \right) \\
 \Rightarrow & \det \left( A \right) = 0.5 + 1.5 = 2 \\ 
\end{gathered} \"


Then 


"{A^{ - 1}} = \frac{{AdjA}}{{\det \left( A \right)}}\"


"{A^{ - 1}} = \frac{\begin{bmatrix}
 0.5 & -3 & 3\\
 0.5 & 1 & -1\\
 -0.25 & 1.5&0.5\\
\end{bmatrix}}{{(2)}}\"



"{A^{ - 1}} = \begin{bmatrix}
 0.25 & -1.5 & 1.5\\
 0.25 & 0.5 & -0.5\\
 -0.125 & 0.75&0.25\\
\end{bmatrix}"


Therefore,


"X=A^{-1}B"


"X=" "\begin{bmatrix}
 0.25 & -1.5 & 1.5\\
 0.25 & 0.5 & -0.5\\
 -0.125 & 0.75&0.25\\
\end{bmatrix}\begin{bmatrix}
 x \\
 y \\
 z\\
\end{bmatrix}"




"\begin{bmatrix}
 x \\
 y \\
 z\\
\end{bmatrix}=" "\begin{bmatrix}
 4\\
 0\\
 1\\
\end{bmatrix}"


Hence


"x=4,\ y=0\ z=1"